OpenStudy (anonymous):

Solve the system of equations by substitution. What is the solution for x? 2x + y = 1 4x + 2y = −1 x = 0 x = 2 There is no x value as there is no solution. x can be any value as there is an infinite number of solutions.

OpenStudy (anonymous):

it is not the same equation

OpenStudy (anonymous):

So Its' There is no x value as there is no solution.

OpenStudy (amistre64):

plug in 0 and then 2 to see if y has a solution

OpenStudy (anonymous):

y=-2x+1 4x+2(-2x+1)=-1 4x-4x+2=-1 0=2 this is not possible, you are correct

Parth (parthkohli):

It's there is no value of x. How can the same number equal to two different numbers?

OpenStudy (anonymous):

2x + y = 1 4x + 2y = −1 rearrange the first like this: y = 1 -2x now sub in this as y in the second: 4x + 2(1-2x) = -1 4x - 4x + 2 = -1 2 = -1 hence inconsistent. no solutions

OpenStudy (anonymous):

also we can see that the equations are parallel, so they would not give a solution?

OpenStudy (anonymous):

they are not parallel

OpenStudy (anonymous):

oh wait, they are, sorry :P

OpenStudy (anonymous):

they are: 2x + y = 1 is the same as y = 1 - 2x 4x + 2y = −1 is the same as y = -0.5 - 2x

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

if they weren't then there would be a solution